OpenStudy (anonymous):

Find the equation in standard form which passes through (-2,-7) and (1,-1). You must show all your work.

OpenStudy (anonymous):

m=(y2-y1)/(x2-x1) to find slope then plug things in (y-y1)=m(x-x1)

OpenStudy (anonymous):

would the answer be 2x - y = -3?

OpenStudy (anonymous):

check your signs pls

OpenStudy (anonymous):

would it be -2x - y =-3..?

OpenStudy (anonymous):

m=(-1+7)/(1+2)=6/3=2 y+1=2(x-1) y+1=2x-2 y=2x-3 -2x+y=-3